OpenStudy (anonymous):

Divide X^(3/5) / X^(1/4)

OpenStudy (angela210793):

=X^(3/5-1/4)=X^(7/20)

OpenStudy (anonymous):

thxs:)

OpenStudy (angela210793):

yw